OpenStudy (zenmo):

Find a unit vector in the direction of the given vector. Verify that the result has a magnitude of 1. W=4j

OpenStudy (anonymous):

need help

jimthompson5910 (jim_thompson5910):

W = 4j = 0i + 4j Anything of the form W = ai + bj has the magnitude |W| = sqrt(a^2 + b^2) The unit vector of W will be (1/|W|)*W this vector will be of length 1 and it will point in the same direction as W

OpenStudy (anonymous):

thats kind of hard

OpenStudy (zenmo):

I still don't understand. :/

jimthompson5910 (jim_thompson5910):

do you see how 0i + 4j matches up with ai + bj making a = 0 b = 4

OpenStudy (zenmo):